(1) if P1 and P2 are prime ideals of a ring R , then P1P2 = P1 intersection P2.
(2) if k is a field , so is k × k
(3) Q[x]/<x^6 + 17 > is a field of characteristic 6.
True/false. Justify.
1
Expert's answer
2020-03-27T14:24:17-0400

1.False.

Let "P_1 \ and \ P_2" are two ideals of a ring "R" .

Then , "P_1\cap P_2=\{ x\in R:x\in P_1 \ and \ x\in P_2 \}"

and "P_1P_2=\{a_1b_1+a_2b_2+.....a_nb_n:a_i\in P_1,b_i\in P_2," "n" is a positive integer "\}" .

Consider the ring "R=\Z"

Let "P_1=2\Z" ,"P_2=2\Z"

Then,"P_1\cap P_2=2\Z" .

Claim:"P_1.P_2=4\Z"

Let "x\in 4\Z. \ Then \ x=4a \ for \ some\ a\in \Z."

Therefore,"x=2a+2a\in P_1.P_2"

"\implies \ 4\Z \subset P_1.P_2"

Let "x\in P_1.P_2."

Then "x=2a_1.2b_1+2a_2.2b_2+.........+2a_r2.b_r \"

for some ,"a_i,b_i \in \Z" .

"\implies\ x= 4(a_1b_1+a_2b_2+.....+a_rb_r)\in 4\Z" .

"\implies\ P_1.P_2\subset 4\Z" .

Hence,"4\Z=P_1P_2."

Therefore,"P_1P_2\neq P_1 \cap P_2."


2.False.

"K×K=\{ (a,b):a,b\in K \}"

Addition and multiplication are defined as follows.

"(a_1,b_1)+(a_2,b_2)=(a_1+a_2,b_1+b_2)."

"(a_1,b_1)(a_2,b_2)=(a_1a_2,b_1b_2)" .

Clearly, "(1,1)" is the multiplicative identity.

Let "(a_1,b_1)\neq(0,0)\in K×K" .

Then "(a_1,b_1)^{-1}" does not exist in general.

for example, "(1,0)\neq(0,0)."

But "(1,0)^{-1}" does not exist.



3.False.

Every field is integral domain by definition and we know that the characteristic of an integral domain is zero or a prime number.

Hence ,6 is not a characteristics of the given field

"\frac{Q[x]}{<x^6+17>}"